# Large Language Model for Qualitative Research - A Systematic Mapping Study

### v1

@@ V Threats to Validity
Construct Validity: A critical issue involves whether the tools and methods used truly capture the concepts under investigation. To ensure that the prompts used were aligned with the research objectives, we adopted an iterative prompt engineering process in which different versions were tested and refined. These versions were made available on Zenodo999https://doi.org/10.5281/zenodo.14177036 to promote greater transparency. Additionally, test articles with known responses were used to validate whether the model’s outputs were consistent with expectations. This procedure provided greater confidence that ChatGPT correctly interpreted the questions and returned responses aligned with the study’s objectives.
@@ VI Conclusion
As future work, we plan to develop a method using LLMs to automate the qualitative analysis process, particularly focusing on the coding stages, by exploring new architectures applied in different tools. For this new method, we also aim to incorporate improvements in prompt engineering techniques to achieve greater accuracy and robustness.
@@ Acknowledgment
The authors express their gratitude for the support of ChatGPT 4.0 (paid version), which was used in the data extraction stage for this systematic mapping. The model was essential in accelerating the process of identifying and selecting relevant articles. Additionally, ChatGPT was also employed for translating texts into English, followed by manual review to ensure accuracy and clarity. All translations and extractions were subsequently verified by human reviewers to ensure spelling, grammar, and consistency of the analyzed data.

### v2

@@ V Threats to Validity
Another limitation concerns the lack of combination with other techniques in the systematic mapping, such as snowballing. Including this technique could contribute to identifying additional relevant studies. According to Wohlin et al. [21], when snowballing is combined with database searches, the review shows a significant increase in the identification of primary studies, allowing gaps left by strategies based solely on database searches to be filled. The absence of this approach may have led to the oversight of relevant articles and represents an opportunity to improve the study.
@@
The decision to exclude studies that did not describe prompt engineering (DE11) reflects the critical need to ensure replicability in this systematic mapping. Given that the use of LLMs for qualitative analysis is a relatively new area of study, detailed descriptions of methodologies, including prompt engineering, are crucial for enabling comparisons and understanding results across studies. However, we recognize that this criterion may have restricted the number of included studies, particularly during the early stages of research in this field. Future iterations of this study could consider less restrictive criteria during the screening phase, enabling a broader analysis that includes studies with incomplete methodological details. This approach could provide valuable insights into the evolution of LLM applications for qualitative analysis.
@@
Construct Validity: A critical issue involves whether the tools and methods used truly capture the concepts under investigation. To ensure that the prompts used were aligned with the research objectives, we adopted an iterative prompt engineering process in which different versions were tested and refined. These versions were made available on Zenodo999https://doi.org/10.5281/zenodo.14177022 to promote greater transparency. Additionally, test articles with known responses were used to validate whether the model’s outputs were consistent with expectations. This procedure provided greater confidence that ChatGPT correctly interpreted the questions and returned responses aligned with the study’s objectives.
@@ VI Conclusion
As future work, we plan to develop a method using LLMs to automate the qualitative analysis process, particularly focusing on the coding stages, by exploring new architectures applied in different tools. For this new method, we also aim to incorporate improvements in prompt engineering techniques to achieve greater accuracy and robustness. The integration of LLMs into structured qualitative analysis workflows also presents an interesting direction for future research, as it may enhance coder agreement and decision-making in collaborative settings.
@@
VII Take Away Lessons
The systematic mapping study provides valuable insights into the application of LLMs for qualitative analysis. Based on the evidence from the analyzed articles, several key takeaways emerge:
VII-A Most Effective LLM According to the Evidence
VII-B Capabilities and Current Limitations
Open Coding: LLMs like ChatGPT are effective in automating the open coding process, rapidly identifying themes and patterns in textual data [14],[8].
Axial Coding and Visualization: Tasks like axial coding and generating graphical representations (e.g., mind maps and visual models in Atlas.TI) still require user intervention. LLMs cannot yet perform complex integrations or visualizations that demand contextual and nuanced decision-making.
Efficiency Gains: A key finding highlights LLMs’ potential to drastically reduce qualitative coding time. For instance, Mathis et al. [9] reported that LLMs shortened coding from weeks to hours, demonstrating their transformative impact on qualitative research workflows.
VII-C Best Practices and Future Directions
While LLMs offer substantial benefits, their effective use requires structured prompts and careful oversight. Researchers should leverage LLMs for automation-friendly tasks, such as initial coding and theme identification, while reserving interpretative or creative processes for human analysts. Advances in LLM architectures and prompt engineering may further expand their role in complex stages of qualitative analysis.
@@ Acknowledgment
The authors thank ChatGPT 4.0 (paid version) for its support in data extraction for this systematic mapping. The model significantly accelerated the identification and selection of relevant articles. It was also used for text translation into English, followed by manual review to ensure accuracy and clarity. All translations and extractions were later verified by human reviewers for spelling, grammar, and data consistency.